Operations Report – January 5th

Supplemental Operations Report 3 5JAN2022

Name of person filing report: Shannon Rupert

Reason for Report: Routine

Non-nominal systems: Nothing to report

Action taken for non-nominal systems: n/a

Generator: Total run time 283.4 hours. We changed the oil today and reset the maintenance warning. Removed all sand from the interior. Still being manually run at night.

ScienceDome Dual Split: Off

Solar—Nominal, charge controllers off at night. We are turning the generator on before the SOC reaches 70% as much as possible.

Solar— VDC Last daylight hours:

Average 51.11

Minimum 46.60

Maximum 59.20

Notes on solar: Nothing to report

Propane Reading, station tank – 45 %

Propane Reading, director tank— 49 %

Propane Reading, intern tank— 50 %

Propane Reading, generator— not checked %

Ethanol Free Gasoline – 2 gallons

Water (Outpost tank) – 250 gallons

Perseverance rover used: yes

Hours: not checked

Beginning charge: 100

Ending charge: not checked

Currently charging: no

Sojourner rover used: yes

Hours: not checked

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Notes on rovers: The dump stabilizer bar on Percy broke and I have ordered a replacement. We aren’t sure if it broke because of the accident, so it should be looked over carefully once it is back in service for any additional breakage the first few times it is used.

ATV’s Used: (Honda, 350.1, 350.2, 300): none

Reason for use: n/a

Oil Added? No

ATV Fuel Used: 0 gallons

# Hours the ATVs were Used: 0

Notes on ATVs: 300 moved into place at the Outpost

HabCar used and why, where? Yes, to town for water and packages

CrewCar used and why, where? Yes, to town for errands

Luna used and why, where? Yes, to town for errands

Business internet: Nothing to report

General notes and comments: Nothing to report.

Campus wide inspection, if action taken, what and why? Nothing to report

Summary of general operations: We cemented the two signs on Cow Dung Road into the ground so they would not blow over anymore. We attempted to replace the toilet in the Hab but had very little time to complete the work and the bolts that secure the toilet to the floor were stuck and we did not have time to unstick them before the crew asked if they could just use the broken one until the end of their rotation. We removed the old diesel overflow berm and releveled the area.
Summary of Outpost operations: Shed was cleaned and organized. Outflow on the water tank was replaced as it had been stripped. Intern Trailer was leveled. Greywater tank control appears to be failing on the intern trailer and may need to be replaced.

Summary of health and safety issues: Nothing to report

Copyright © The Mars Society. All rights reserved. | Main Site